### v2

# I/O in Machine Learning Applications on HPC Systems: A 360-degree Survey

Abstract.
Growing interest in Artificial Intelligence (AI) has resulted in a surge in demand for faster methods of Machine Learning (ML) model training and inference. This demand for speed has prompted the use of high performance computing (HPC) systems that excel in managing distributed workloads. Because data is the main fuel for AI applications, the performance of the storage and I/O subsystem of HPC systems is critical. In the past, HPC applications accessed large portions of data written by simulations or experiments or ingested data for visualizations or analysis tasks. ML workloads perform small reads spread across a large number of random files. This shift of I/O access patterns poses several challenges to modern parallel storage systems. In this paper, we survey I/O in ML applications on HPC systems, and target literature within a $6$-year time window from $2019$ to $2024$. We define the scope of the survey, provide an overview of the common phases of ML, review available profilers and benchmarks, examine the I/O patterns encountered during offline data preparation, training, and inference, and explore I/O optimizations utilized in modern ML frameworks and proposed in recent literature. Lastly, we seek to expose research gaps that could spawn further R&D.
I/O access pattern, HPC I/O, storage, machine learning
journal: CSURjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3722215ccs: Information systems Information storage systemsccs: Information systems Storage architecturesccs: Information systems Hierarchical storage managementccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning
1. Introduction
Existing High-Performance Computing (HPC) Input and Output (I/O) research has focused on techniques to optimize performance when running traditional HPC application workloads, which typically include simulations and checkpointing partial results (et al, 2021e). Because of the increased popularity of Machine Learning (ML) workloads, there is a rising demand for I/O systems that can effectively accommodate their distinct I/O access patterns. Write operation bursts commonly dominate traditional workloads; however, ML workloads are usually read-intensive and use many small files (et al, 2020f). Due to the absence of a well-established consensus on the preferred I/O stack for ML workloads, numerous developers resort to crafting their own ad hoc algorithms and storage systems to accommodate the specific requirements of their applications (Gainaru et al., 2022). This can result in suboptimal application performance due to the under-utilization of the storage system, prompting the necessity for novel I/O optimization methods tailored to the demands of ML workloads.
Figure 1. Traditional HPC I/O stack vs evolving ML I/O HPC stack (adapted from (et al, 2021e)).
In Fig. 1, we show the I/O stack used to run ML workloads (on the right side) compared to the traditional HPC I/O stack (on the left side) (et al, 2021e). The traditional HPC I/O stack has been developed to support massive parallelism. Applications are typically written in C/C++/Fortran languages that use the MPI programming model and I/O middleware, such as HDF5 (The HDF Group, 2022; Byna et al., 2020) and netCDF (Lee et al., 2008), which internally use MPI-IO and POSIX-IO interfaces to move (read or write) data between main memory and parallel file systems (PFS). In contrast, ML workloads are typically written in productivity-oriented languages, such as Python, and use PyTorch and TensorFlow frameworks.
Although many ML workloads still use the traditional I/O stack on HPC systems, some are shifting to use remote cloud computing systems that offer services such as data warehouses to store and retrieve large datasets. Cloud platforms such as Google Cloud Platform, AWS, and Microsoft Azure offer access to large datasets, including Google Analytics data (Google Cloud, 2024a) hosted on BigQuery (Google Cloud, 2024b), the $1000$ Genomes Project (Amazon Web Services, 2024a) hosted on Amazon Redshift (Amazon Web Services, 2024b), and NYC Taxi & Limousine Commission data (Microsoft Azure, 2024b) hosted on Azure Synapse Analytics (Microsoft Azure, 2024a). HPC workloads that aim to use these datasets require near-first-class integration with the cloud to enable efficient data transfers to HPC systems. When using the traditional I/O stack, high bandwidth and low latency data access are possible due to the locality of the data and the highly optimized stack. In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding and Taylor, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024). Given the high latency inherent in accessing data stored in a remote storage location, optimizing data retrieval is crucial for reasonable performance. Despite the lower I/O performance imposed when using a data warehouse, they are still preferred in many scenarios because they provide a centralized data storage location, facilitate data integration from various sources, and enable data governance, which can result in higher-quality data. High-quality data is essential when training ML models because it directly influences their accuracy, reliability, and generalizability (et al, 2020c). In addition, the monetary cost needed to purchase and maintain an HPC system can far outweigh the cost of using managed cloud services (Carlyle et al., 2010; et al, 2013). Furthermore, integration of HPC systems with available cloud resources is necessary due to the cloud being used as a data storage solution from which datasets can be publicly shared and used by all.
The differences listed above in the I/O stacks for ML applications using the traditional PFS compared to cloud computing systems motivate the development of novel storage system designs and I/O performance optimization techniques to adapt to the I/O operations commonly found in ML processes. Ideally, the methods would be transparent to the application developers. Notably, many scientists may lack formal training in software development, as evidenced by Paul et al. (et al, 2021g). The study revealed that ML workloads are becoming ever more popular and that developers from various scientific domains beyond computer science often overlook the utilization of node-local storage (i.e., burst buffers) within HPC systems. Further evidence for the increasing interest in distributed training is the number of posts on the popular tech website StackOverflow. The cumulative number of posts discussing distributed training increased staggeringly from $257$ posts in $2015$ to $2,061$ posts in 2024 (Community, 2024). The increasing number of ML workloads that are running on HPC systems prompts the need for transparent I/O optimization techniques to decrease the complexity required for scientists to write software that takes advantage of the resources offered by HPC systems.
Furthermore, scientists, I/O performance enthusiasts, and HPC practitioners can gain significant insights from benchmarks that accurately emulate ML workloads and profilers that capture critical I/O information from the underlying storage systems. Using these tools, individuals can find potential I/O bottlenecks within their ML applications. Access to realistic benchmarking tools and fine-grained profilers is also invaluable for HPC storage research, allowing researchers to analyze and emulate the diverse range of I/O operations found in ML workloads. Currently, I/O analysis tools often fall short in a number of ways, including the inability to identify I/O requests at the thread level and only capturing traces of specific I/O interfaces (i.e., POSIX and STDIO). This makes it challenging to find a general-purpose tool applicable to a wide variety of workloads, which can reduce the consistency in which I/O traces are captured. In addition, many tools (i.e., Darshan (et al, 2016c) and Scalasca (Geimer et al., 2010)) target MPI applications limiting their compatibility and effectiveness with other communication libraries (i.e., GLOO (Facebook Incubator, 2024), Spark (Apache Software Foundation, 2024a), and Dask (Dask, 2024)). It should be noted that Darshan is capable of capturing I/O traces from applications that do not use MPI through the NONMPI environment variable. However, this results in the output of a single DXT log for each process instead of the more common single file output that captures all I/O traces from an application using MPI. In some cases, this leaves developers with the task of manually merging the results from their non-MPI applications in order to conduct further analysis.
The topic of ML I/O and whether it has a significant role in poor performance has been a debate within the HPC community. Many point out that HPC systems often have enough node-local storage to cache the dataset, resulting in minimal time spent waiting on I/O (Schimmelpfennig et al., 2021; et al, 2021h, g, 2018d). Others mention that not all HPC systems have large amounts of node local storage or the datasets used are simply too large to be cached (Schimmelpfennig et al., 2021; et al, 2021h; Xu et al., 2021; et al, 2018d). In addition, due to the random sampling that commonly occurs during ML training, modern PFSs can often become an I/O bottleneck (Petersen, 2015; et al, 2020k, 2018d). Ultimately, whether I/O is a significant issue is subjective to the workload at hand. It depends on various factors such as dataset size, storage capacity, storage system bandwidth, I/O library implementation and their configurations, and I/O access patterns of applications. Furthermore, many ML application developers are not HPC storage experts; hence, even if their system is capable of providing great performance, transparent solutions such as caching and prefetching are needed to ensure the HPC systems are utilized effectively. Transparent solutions are of particular interest due to the increasing number of heterogeneous processors used within ML applications. These increase the complexity of optimally utilizing available compute resources. For these reasons, we believe that I/O and data management are key areas of interest to the ML community.
There are several common phases in an ML application lifecycle: data generation, data preparation, training, and inference. The data management found within each phase varies widely due to different I/O access patterns, data sources, and file formats. We provide a taxonomy of the data management found throughout the ML lifecycle in Fig. 3. The ML lifecycle begins with data generation, which involves collecting data from various sources (e.g., simulation results or web scraping), which are then stored using a chosen file format (e.g., .tfrecord, .csv, or .png) and storage location (e.g., SSD, disk, cloud, or database). This is commonly followed by an offline data preparation phase, which performs a series of transformations and reductions (i.e., filters, aggregations, or generalizations) over the collected data to increase the training efficiency and accuracy of the ML model. The training process involves reading samples from the data sources, which are then fed to computations. The order in which the samples are read largely depends on the chosen optimizer (e.g., Stochastic Gradient Descent and Adaptive Moment Estimation or Adam, etc.) and training distribution strategy (e.g., model or data parallelism). Due to many optimizers requiring random sampling and storage systems being optimized for contiguous sequential reading, training I/O is a potential bottleneck. Once the model has been trained, it is often deployed and fed “new” data during the inference phase. It is common for the online and offline data preparation phases to occur before and during the inference phase as well. More details on each phase of ML can be found in Section 3, but it is important to note the wide range of factors that impact I/O and data management throughout the ML lifecycle, such as varying file formats, data sources, data preparation requirements, and data modalities, among others. It is impractical to cover all possible combinations of the factors mentioned above, and therefore, in this paper, we aim to discuss the I/O approaches found in the most common ML applications running on HPC systems.
The structure of the paper is as follows. First, in Section 2, we define the scope of the survey and present trends in the cited literature. Then, in Section 3, we describe the data formats and modalities found in ML workloads and provide an overview of the typical phases of ML applications and the popular methods of distributing the model training on HPC systems. Next, in Section 4, we describe the benchmarks and profiling tools scientists and researchers can use to ensure proper system utilization under ML workloads. In Section 5, we describe the I/O access patterns found in the offline data preparation, training phase, and inference phases. The training phase is exemplified by analyzing the currently available representative I/O benchmarks. We then discuss, in Section 6, various I/O optimization techniques used in modern ML frameworks and those proposed in the recent research literature. Lastly, in Section 7, we discuss the gaps in the current research of ML I/O in HPC systems.
2. Scope of the Survey
ML applications use various environments, from local machines for small-scale applications to HPC and cloud systems for large-scale modeling. This survey aims to provide a comprehensive review of the I/O patterns in ML running on HPC systems. This means that I/O patterns and other data management solutions that may be common in other environments (e.g., cloud or local) are not relevant. For example, in Sec 5, we discuss only the I/O access patterns found in common HPC scenarios and exclude I/O access patterns that may be common in cloud or local environments.
GPU Mem. Hierarchy
File Format
Sample Caching
Sample Prefetching
Sample Shuffling
Asynchronous Sampling
Profiling
Benchmarking
Model Checkpointing
Model Storage Offloading
Scheduling
ML Analysis
Papers
Acknowledgements.
This research was supported by the U.S. Department of Energy (DOE), Office of Science, Office of Advanced Scientific Computing Research (ASCR) under contract number DE-AC02-05CH11231 with LBNL. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231.

### v3

@@
 When using the traditional I/O stack, high bandwidth and low latency data access are possible due to the locality of the data and the highly optimized stack.
-In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding and Taylor, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024).
+In contrast, access to data stored in a data warehouse is often accomplished through a network interface such as REST (Fielding, 2000), SOAP (Box et al., 2000), or ODBC (Microsoft, 2024).
 Given the high latency inherent in accessing data stored in a remote storage location, optimizing data retrieval is crucial for reasonable performance.
@@
 The topic of ML I/O and whether it has a significant role in poor performance has been a debate within the HPC community.
-Many point out that HPC systems often have enough node-local storage to cache the dataset, resulting in minimal time spent waiting on I/O (Schimmelpfennig et al., 2021; et al, 2021h, g, 2018d).
-Others mention that not all HPC systems have large amounts of node local storage or the datasets used are simply too large to be cached (Schimmelpfennig et al., 2021; et al, 2021h; Xu et al., 2021; et al, 2018d).
-In addition, due to the random sampling that commonly occurs during ML training, modern PFSs can often become an I/O bottleneck (Petersen, 2015; et al, 2020k, 2018d).
+Many point out that HPC systems often have enough node-local storage to cache the dataset, resulting in minimal time spent waiting on I/O (Schimmelpfennig et al., 2021; Farrell et al., 2021; et al, 2021g, 2018c).
+Others mention that not all HPC systems have large amounts of node local storage or the datasets used are simply too large to be cached (Schimmelpfennig et al., 2021; Farrell et al., 2021; Xu et al., 2021; et al, 2018c).
+In addition, due to the random sampling that commonly occurs during ML training, modern PFSs can often become an I/O bottleneck (Petersen, 2015; et al, 2020i, 2018c).
 Ultimately, whether I/O is a significant issue is subjective to the workload at hand.
@@
 ML Analysis
-Papers
+apers
 Acknowledgements.